OpenStudy (gabylovesyou):

If the sin 90° is 1, then the cos ____ °= _____. 90° ; 1 over 2 0° ; square root 2 over 2 90° ; square root 3 over 2 0° ; 1

OpenStudy (gabylovesyou):

@rational

OpenStudy (anonymous):

im confuzzled :D

OpenStudy (anonymous):

Cos0 is 1

OpenStudy (rational):

hint : \( \sin (90 - \theta ) = \cos (\theta)\)

OpenStudy (rational):

when \(\theta =0\) : \( \sin (90 ) = 1 = \cos (0)\)

OpenStudy (gabylovesyou):

thank you

OpenStudy (rational):

you're welcome !